OpenStudy (anonymous):

olve. The average value of a certain type of automobile was $13,320 in 1995 and depreciated to $4020 in 2000. Let y be the average value of the automobile in the year x, where x = 0 represents 1995. Write a linear equation that models the value of the automobile in terms of the year x. A) y = -1860x + 4020 B) y = - x - 4020 C) y = -1860x + 13,320 D) y = -1860x - 5280

hero (hero):

This is just a simple y = mx + b problem

OpenStudy (anonymous):

k thankS

OpenStudy (usukidoll):

d i s t a n c...

OpenStudy (usukidoll):

wait may be different on this one. ummm..slope formula.

hero (hero):

You mean y = mx + b

hero (hero):

I'm going to plug the points into y = mx + b

OpenStudy (usukidoll):

ya but gotta find my m : )

OpenStudy (usukidoll):



hero (hero):

Well, actually, x = 0 was supposed to be 1995

hero (hero):

But at least we have the slope

hero (hero):

I was supposed to use (0, 13320) and (5, 4020) as the points

OpenStudy (anonymous):

yeah so the equation is?

OpenStudy (anonymous):

i think its c

hero (hero):

y = mx + b 13320 = 0m + b 4020 = 5m + b Isolate b in both to get b = 13320 b = 4020 - 5m 13320 = 4020 - 5m 13320 - 4020 = -5m 9300 = -5m 9300/-5 = m -1860 = m y = 1860x + 13320

hero (hero):

y = -1860x + 13320

OpenStudy (anonymous):

haha i was right!

hero (hero):

Yes, but I showed the work. You should appreciate my methods. My methods are direct application of basic algebraic formulas

hero (hero):

I don't need to use silly distance formula for slope formula which are really sideline formulas.

OpenStudy (anonymous):

yes i had learn that method to but i also went with my instincts lol

hero (hero):

y = mx + b and a^2 + b^2 = c^2 are the main formulas

OpenStudy (anonymous):

thank you! yes i learned the slope method to.

hero (hero):

I bet even @UsukiDoll learned something here. Probably never seen a problem solved in this manner before.

OpenStudy (usukidoll):

This problem doesn't exist in Calculus deary.

hero (hero):

lol, of course it doesn't because it is an Algebra problem.

hero (hero):

I'm familiar with Calculus to btw

hero (hero):

My strange methods are also used there as well

OpenStudy (anonymous):

so for the mx+b this question ask Use the given conditions to write an equation for the line in slope-intercept form. Slope = 2, passing through (-8, 8) A) y - 8 = x + 8 B) y = 2x + 24 C) y = 2x - 24 D) y - 8 = 2x + 8 So do i work it all the way out thus getting c or do i stop at d

hero (hero):

Actually, all you have to do is just insert the points into the formula y = mx + b

OpenStudy (usukidoll):

sorry I gotta practice my own probs before I go to bed.

OpenStudy (anonymous):

I did so its D.

hero (hero):

So let's try it y = mx + b 8 = 2(-8) + b 8 = -16 + b 8 + 16 = b 24 = b y = 2x +24

OpenStudy (anonymous):

I thought u dont work it all the way out

hero (hero):

Why wouldn't you?

hero (hero):

We had everything. All we had to do is find b. Notice that my result is one of the answer choices.

hero (hero):

As I always say, all you need is y = mx + b

OpenStudy (anonymous):

oh ok slope interc. thank you yes that was one of my choice i was just confused.